# The Tantalizing Case of the Quasar J09505128 – I. Presentation of the Data and Detailed Exploration of the Binary Supermassive Black Hole Scenario

### v3

@@ 3.5 The Radial Velocity Curve
Note. — Section 3.5 describes how the absolute shifts were obtained. In addition to the tabulated uncertainties, there is a systematic uncertainty of km s-1 from Fe ii subtraction (Section 3.2.2), a systematic uncertainty of km s-1 due to the increase in the width of the broad H profile with time (Section 3.3), and a 200 km s-1 jitter (Section 5.3).
@@
3.6 A Summary of Uncertainties
Throughout Section 3, we measured velocity shifts of the broad H line and did a series of tests and simulations to characterize the uncertainties in those measurements. The formal uncertainties from cross correlation are described in Section 3.2 and are listed in Table 3; they are between 40 and 60 km s-1. A systematic uncertainty of about 60 km s-1 arises from the modeling and subtraction of the Fe ii complex (Section 3.2.2). From the simulation in Section 3.3, we estimated an uncertainty of 80 km s-1 due to the increasing width of the broad H line over time. To account for short-timescale radial velocity fluctuations, such as the example in Section 3.2.1), we adopted a jitter of 200 km s-1 (discussed further in Section 5.3) for use in the radial velocity analyses presented later. The jitter is the dominant source of uncertainty.
4 Analysis of Light Curves
@@
In addition to the periodograms, we modeled each light curve with a flat line corresponding to its 3-clipped mean magnitude, and computed residuals normalized by the measurement uncertainties, as shown in Figure 11. The distributions of residuals normalized by the error bars appear approximately symmetric, though with a broader spread than a standard normal distribution, likely due to intrinsic stochastic variability of the quasar (e.g., vaughan2016false; liu2018did). The fit of a flat line in Figure 11 and the lack of any any periodic signal in the periodograms, leads us to conclude that there is no evidence of periodicity in the light curves. The light curves of J0950 are revisited in Section 5.5.
5 Interpretation, Discussion, and Speculation
5.1 Summary of Main Results from the Data
@@
5.2 Assessment of Different Interpretations for the Observed Velocity Shifts
In this section, we consider the interpretations listed in Section 1.2 for the observed shift of the broad H line of J0950. We discuss those interpretations and evaluate whether they are physically plausible and applicable, before contemplating whether they can be subjected to statistical tests.
The broad H velocity shifts in the spectra of J0950 have increased steadily over the decades it has been monitored (Section 3.5). Such a trend resembles closely the expected signature of a long-period (several decades) binary supermassive black hole with a BLR associated with one of the two objects. The shifts in the broad H are then associated with the changes of the projected bulk velocity of the BLR over a portion of an orbital cycle. Since the time scale and general pattern of the radial velocity variations is consistent with expectations (e.g., eracleous2012large; runnoe2017large), we explore this hypothesis further by fitting a Keplerian orbit model to the observed radial velocity curve in Section 5.3 and considering the implications in Sections 5.4 and 5.5.
@@ 5.2 Assessment of Different Interpretations for the Observed Velocity Shifts
We can also estimate the crossing time for an obscuring object orbiting the BLR by following the method of lamassa2015. For a supermassive BH with a BLR radius of pc (see estimates in Section 5.4), it would take years for a dust cloud in Keplerian orbit at that radius to cross the BLR. Given that the broad H line of J0950 exhibits a significant blueshift and redshift in the span of years, this timescale discrepancy makes it difficult to reconcile the observations with the dust-obscuration scenario. Furthermore, matching the observed timescale of about 20 years would require a BH mass , which is physically implausible (e.g., thomas201617; inayoshi2016there; mcconnell2012dynamical).
@@
Another possible source of shifted broad emission lines is outflows. In the case of bipolar outflows (e.g., zheng1990outflows; veilleux1991outflows; zheng1991outflows) where the ejection is symmetric and expands outwards, the observer would first see the blueshifted side, followed by the redshifted side after a time delay. In this scenario, both the red and blue sides of a broad emission line must be visible. The only way for the blue side to appear first, then disappear, and then the red side appear, as we observe in J0950, would be if the blue side were to turn off exactly as the red side turns on. This would require exquisite fine tuning. Even with such fine tuning, the bipolar outflow scenario cannot explain the symmetric 2015 line profile centered at .
@@
The possibility that J0950 is a recoiling BH is unlikely on several grounds. A recoiling BH, after being ejected, could reach a maximum distance from the galactic center up to several kpc (e.g., blecha2016recoil) before falling back and oscillating around the nucleus. However, the timescale for one full oscillation cycle is at least years (komossa&meritt2008recoil), which is far longer than the duration over which J0950 has been observed. Furthermore, although recoil speeds in excess of several thousand are possible in principle (e.g., healy2023recoil), the probability of observing a kick velocity that is from a supermassive binary merger is negligibly small; most recoil velocities are on the order of a few hundred or less (dotti2010dual).777Recoils from three-body interactions, can reach 1000 km s-1 for the lightest supermassive BH in a three-body interaction but are typically on the order of a few hundred (hoffman2006recoil). The likelihood of detecting kick velocities much greater than the escape speed of the host galaxy is even lower. In the case of J0950, where we observe velocities as high as , that probability is negligibly small (blecha2016recoil). Most critically, the radial velocity curve of J0950 mostly shows consistent positive acceleration as well as a change in sign––a behavior that is inconsistent with the expected dynamics of a recoiling BH.
@@
There are known objects that show a single displaced peak with a shoulder, similar to the earliest spectrum of J0950. Examples include Mrk 668 and 3C 227, both of which have broad emission lines with a displaced peak and shoulder (marziani1993peculiar; gezari2007long). They have been monitored, more sparsely than J0950, and some do show radial velocity variations, but none have the pattern we see in J0950 (especially not the change of sign of the radial velocity curve). Non-axisymmetric disk models (e.g., storchi2017double; schimoia2017evolution) have been used to try to explain their irregular, non-periodic radial velocity curves. These models entail a “perturbation”, a bright spot or spiral arm, orbiting in the inner region of a circular, disk-like BLR, or a precessing elliptical disk (gezari2007long; lewis2010long; storchi2017double). This scenario can cause a broad emission line with two peaks, whose strengths alternate like a seesaw as the perturbation moves and enhances one peak after another. The smoothly increasing radial velocity variation of J0950 differs from the variations of these objects and cannot be explained by the non-axisymmetric BLR models described above. Nonetheless, the broader framework of a perturbed BLR may provide a plausible explanation for J0950’s behavior. Since other objects within this framework have shown substantial changes in their radial velocity over time, continued monitoring of J0950 is essential to see whether its trend persists or evolves (see the discussion of 3C 390.3 in Section 4.2 of eracleous1997rejection). Further testing of the single perturbed BLR scenario for J0950 will require developing a detailed physical model, which is out of the scope for this paper.
@@
J0950 is an unusual object. Among quasars, only double-peaked emitters show behavior even remotely resembling the large monotonic velocity variations seen in J0950. doan2020improved present the longest monitoring campaign for double-peaked emitters; only two objects (3C 332 and 3C 390.3) displayed fairly monotonic velocity changes over most of their monitoring periods. Since double-peaked emitters make up only 3% of all quasars (strateva2003doublepeaked), such large monotonic variations are extremely rare––roughly 0.4% of quasars (2/14 of extensively monitored double-peaked emitters 3%). Importantly, none of these objects show a change in sign in their radial velocity curves. Therefore, the velocity changes observed in J0950 cannot be directly attributed to phenomena seen in other quasars.
@@
5.3 A Supermassive Binary Black Hole Scenario for the Radial Velocity Curve
@@
The radial velocity curve of J0950 based on the fourth column of Table 3 is shown in Figure 12. It shows a steady increase of 3500 km s-1 over time; the most striking behavior is that the broad line region of this quasar has undergone radial velocity variations in a manner resembling Keplerian motion in a binary. There certainly are potential caveats, which we discuss in Section 5.2. But the radial velocity curve of J0950 so far is consistent with binary motion, which makes it a strong SBHB candidate.
@@
We increased the measurement uncertainties of the radial velocity curve to capture the effects of “jitter” from regular quasar variability (guo2019constraining; doan2020improved). Quasars can have perturbations or asymmetries within the broad line region that manifest as small, stochastic radial velocity variations, or jitter, occurring over a timescale of the order of a year or less (e.g., barth2015lick). This subtle effect is directly measurable in two of J0950’s spectra (see Section 3.2.1). doan2020improved empirically determined that jitter can be characterized as fluctuations around a smooth radial velocity curve, distributed approximately according to a Gaussian distribution (see also guo2019constraining). The standard deviation of the distribution is on the scale of a few hundred km s-1. Following the methodology from that work, we added a jitter of 200 km s-1 in quadrature to the uncertainties in each radial velocity measurement. The jitter dominates over the measurement errors and is the primary source of uncertainty in the radial velocity at each epoch.
@@
To fit the radial velocity curve shown in Figure 12, we used radvel (fulton2018radvel), a package for fitting Keplerian orbits to radial velocity time series. The best-fitting orbital parameters, determined from the radial velocity curve fit in the observed frame, are provided in Table 4. The free parameters in the fitting are the period (), the time of conjunction (), the natural log of the semi-amplitude (ln ), and and , from which the eccentricity () and argument of periapsis () are derived. Given the exploratory nature of fitting the velocity curve of J0950 as a binary and the absence of prior knowledge about the system, we adopted broad, physically motivated priors based on the radial velocity curve. Specifically, we assumed a near-circular orbit as a starting point and adopted Gaussian priors on the period and time of conjunction: a mean period of yr (ensuring that the 1- range does not fall below the 22 yr span of radial velocity measurements available), and a time of conjunction yr. For a circular orbit, the conjunction––the moment when the lower-mass, or secondary BH, passes along the line of sight between the observer and the primary BH––would be expected near that epoch. A hard bound of was applied to the semi-amplitude to restrict unphysical amplitudes, and an eccentricity prior of was imposed. The Markov chain Monte Carlo (MCMC) was run with 20 walkers for 2500 steps each, and the chains converged reliably after steps (Gelman-Rubin statistic 1.01) following an initial burn-in. The resulting posteriors were well-constrained and consistent with the observed signal (see Appendix B for a corner plot), indicating that the priors were non-restrictive.
@@
Since the radial velocity baseline of 22 yr covers only a fraction of the 33-yr orbit, the period, time of conjunction and the semi-amplitude, become partially degenerate. Therefore, the posterior distribution at longer periods is strongly influenced by the assumed prior. For much longer periods, the radial velocity curve would look almost the same over the observed 22 yr window. Likewise, the semi-amplitude can shift slightly depending on the adopted jitter value. These effects are expected for long period systems.
@@
5.4 Exploration of the Supermassive Binary Black Hole Scenario and Discussion
@@
A lower limit on the total mass of the SBHB in J0950 was initially estimated at M⊙ by runnoe2017large. We obtained limits on the mass assuming that the bolometric luminosity does not exceed the Eddington luminosity. The bolometric luminosity was estimated by taking the luminosity of J0950’s 2002 spectrum––the brightest state of this object––at 5100 Å, and applying a bolometric correction factor of 10.3 (richards2006spectral). This yields a mass of . We estimated another lower mass limit, taking maximum inclination () and a mass ratio of , in the following Keplerian velocity equation,
@@
This equation was derived under the assumption of an unequal mass binary system, in which the less massive, or secondary, BH is active and its BLR emits the observed H line (see arguments in eracleous2012large; runnoe2015large). Using all combinations explored by the MCMC algorithm for the period , the velocity semi-amplitude , and eccentricity , the inferred 68%-confidence lower limit on total mass is and the 99% lower limit is . These inferred masses do not necessarily correspond to the lower limits on , , and . All the lower mass limits above are reasonable in the sense that they are less than , a threshold regarded as unphysical. This threshold comes about because the most massive BHs measured are of the order of and because BHs are not expected to accrete material rapidly enough to surpass this mass (e.g., thomas201617; inayoshi2016there; mcconnell2012dynamical). In line with this argument, the minimum possible inclination for J0950 is . Using the distributions of period and mass, applying Kepler’s law gives 68% and 99% lower limits on the orbital semi-major axis of 0.008 pc and 0.006 pc respectively.
@@
If J0950 is a binary, we attribute the offset broad H emission lines to a BLR associated with the lower mass, or secondary, BH. In this scenario, the tidal interaction with the companion BH may limit the size of the BLR. To explore that possibility, we estimate the “characteristic radius” of the BLR888This is the length that corresponds to the observed time delay, between fluctuations of continuum and a broad line, . using the empirical relation from bentz2013 between the characteristic radius and AGN luminosity at 5100 Å,
@@
with and . We applied this relation even though it is appropriate only for a single BH and does not account for the dynamical effects in a binary system to get a characteristic radius of the BLR of pc. This value is somewhat higher than our inferred lower limit on the binary semi-major axis of pc. The extent of the line-emitting region in the BLR is possibly 45 times greater than (see discussion in runnoe2015large). Since we only have a lower limit on the orbital semi-major axis, there appears to be no conflict between this separation and the characteristic radius of the BLR.
@@
If J0950 does indeed harbor a binary, extrapolating from the fit to its radial velocity curve, it should start turning over soon. It is therefore critical that we observe this object in the near future, at a rate of at least one spectrum per year to see if this prediction is borne out. It is also important to note that our implementation of radial velocity jitter discussed in Section 3.5 is simple, and more sophisticated prescriptions of jitter should be developed and employed. The description of jitter would impact the uncertainties on the orbital parameters that are inferred from fitting a radial velocity curve. Regardless of whether or not J0950 is a binary, it still displays an exceptional behavior that warrants detailed study.
5.5 On the Detectability of Doppler Boosting in the Light Curves

### v4

@@ III.5 The Radial Velocity Curve
Note. — Section III.5 describes how the absolute shifts were obtained. In addition to the tabulated uncertainties, there is a systematic uncertainty of km s-1 from Fe ii subtraction (Section III.2.2), a systematic uncertainty of km s-1 due to the increase in the width of the broad H profile with time (Section III.3), and a 200 km s-1 jitter (Section V.3).
@@
III.6 A Summary of Uncertainties
Throughout Section III, we measured velocity shifts of the broad H line and did a series of tests and simulations to characterize the uncertainties in those measurements. The formal uncertainties from cross correlation are described in Section III.2 and are listed in Table 3; they are between 40 and 60 km s-1. A systematic uncertainty of about 60 km s-1 arises from the modeling and subtraction of the Fe ii complex (Section III.2.2). From the simulation in Section III.3, we estimated an uncertainty of 80 km s-1 due to the increasing width of the broad H line over time. To account for short-timescale radial velocity fluctuations, such as the example in Section III.2.1), we adopted a jitter of 200 km s-1 (discussed further in Section V.3) for use in the radial velocity analyses presented later. The jitter is the dominant source of uncertainty.
IV Analysis of Light Curves
@@
In addition to the periodograms, we modeled each light curve with a flat line corresponding to its 3-clipped mean magnitude, and computed residuals normalized by the measurement uncertainties, as shown in Figure 11. The distributions of residuals normalized by the error bars appear approximately symmetric, though with a broader spread than a standard normal distribution, likely due to intrinsic stochastic variability of the quasar (e.g., Vaughan et al., 2016; Liu et al., 2018). The fit of a flat line in Figure 11 and the lack of any any periodic signal in the periodograms, leads us to conclude that there is no evidence of periodicity in the light curves. The light curves of J0950 are revisited in Section V.5.
V Interpretation, Discussion, and Speculation
V.1 Summary of Main Results from the Data
@@
V.2 Assessment of Different Interpretations for the Observed Velocity Shifts
In this section, we consider the interpretations listed in Section I.2 for the observed shift of the broad H line of J0950. We discuss those interpretations and evaluate whether they are physically plausible and applicable, before contemplating whether they can be subjected to statistical tests.
The broad H velocity shifts in the spectra of J0950 have increased steadily over the decades it has been monitored (Section III.5). Such a trend resembles closely the expected signature of a long-period (several decades) binary supermassive black hole with a BLR associated with one of the two objects. The shifts in the broad H are then associated with the changes of the projected bulk velocity of the BLR over a portion of an orbital cycle. Since the time scale and general pattern of the radial velocity variations is consistent with expectations (e.g., Eracleous et al., 2012; Runnoe et al., 2017), we explore this hypothesis further by fitting a Keplerian orbit model to the observed radial velocity curve in Section V.3 and considering the implications in Sections V.4 and V.5.
@@ V.2 Assessment of Different Interpretations for the Observed Velocity Shifts
We can also estimate the crossing time for an obscuring object orbiting the BLR by following the method of LaMassa et al. (2015). For a supermassive BH with a BLR radius of pc (see estimates in Section V.4), it would take years for a dust cloud in Keplerian orbit at that radius to cross the BLR. Given that the broad H line of J0950 exhibits a significant blueshift and redshift in the span of years, this timescale discrepancy makes it difficult to reconcile the observations with the dust-obscuration scenario. Furthermore, matching the observed timescale of about 20 years would require a BH mass , which is physically implausible (e.g., Thomas et al., 2016; Inayoshi and Haiman, 2016; McConnell et al., 2012).
@@
Another possible source of shifted broad emission lines is outflows. In the case of bipolar outflows (e.g., Zheng et al., 1990; Veilleux and Zheng, 1991; Zheng et al., 1991) where the ejection is symmetric and expands outwards, the observer would first see the blueshifted side, followed by the redshifted side after a time delay. In this scenario, both the red and blue sides of a broad emission line must be visible. The only way for the blue side to appear first, then disappear, and then the red side appear, as we observe in J0950, would be if the blue side were to turn off exactly as the red side turns on. This would require exquisite fine tuning. Even with such fine tuning, the bipolar outflow scenario cannot explain the symmetric 2015 line profile centered at .
@@
The possibility that J0950 is a recoiling BH is unlikely on several grounds. A recoiling BH, after being ejected, could reach a maximum distance from the galactic center up to several kpc (e.g., Blecha et al., 2016) before falling back and oscillating around the nucleus. However, the timescale for one full oscillation cycle is at least years (Komossa and Merritt, 2008), which is far longer than the duration over which J0950 has been observed. Furthermore, although recoil speeds in excess of several thousand are possible in principle (e.g., Healy and Lousto, 2023), the probability of observing a kick velocity that is from a supermassive binary merger is negligibly small; most recoil velocities are on the order of a few hundred or less (Dotti et al., 2010).777Recoils from three-body interactions, can reach 1000 km s-1 for the lightest supermassive BH in a three-body interaction but are typically on the order of a few hundred (Hoffman and Loeb, 2006). The likelihood of detecting kick velocities much greater than the escape speed of the host galaxy is even lower. In the case of J0950, where we observe velocities as high as , that probability is negligibly small (Blecha et al., 2016). Most critically, the radial velocity curve of J0950 mostly shows consistent positive acceleration as well as a change in sign––a behavior that is inconsistent with the expected dynamics of a recoiling BH.
@@
There are known objects that show a single displaced peak with a shoulder, similar to the earliest spectrum of J0950. Examples include Mrk 668 and 3C 227, both of which have broad emission lines with a displaced peak and shoulder (Marziani et al. 1993; Gezari et al. 2007). They have been monitored, more sparsely than J0950, and some do show radial velocity variations, but none have the pattern we see in J0950 (especially not the change of sign of the radial velocity curve). Non-axisymmetric disk models (e.g., Storchi-Bergmann et al. 2017; Schimoia et al. 2017) have been used to try to explain their irregular, non-periodic radial velocity curves. These models entail a “perturbation”, a bright spot or spiral arm, orbiting in the inner region of a circular, disk-like BLR, or a precessing elliptical disk (Gezari et al., 2007; Lewis et al., 2010; Storchi-Bergmann et al., 2017). This scenario can cause a broad emission line with two peaks, whose strengths alternate like a seesaw as the perturbation moves and enhances one peak after another. The smoothly increasing radial velocity variation of J0950 differs from the variations of these objects and cannot be explained by the non-axisymmetric BLR models described above. Nonetheless, the broader framework of a perturbed BLR may provide a plausible explanation for J0950’s behavior. Since other objects within this framework have shown substantial changes in their radial velocity over time, continued monitoring of J0950 is essential to see whether its trend persists or evolves (see the discussion of 3C 390.3 in Section 4.2 of Eracleous et al. 1997). Further testing of the single perturbed BLR scenario for J0950 will require developing a detailed physical model, which is out of the scope for this paper.
@@
J0950 is an unusual object. Among quasars, only double-peaked emitters show behavior even remotely resembling the large monotonic velocity variations seen in J0950. Doan et al. (2020) present the longest monitoring campaign for double-peaked emitters; only two objects (3C 332 and 3C 390.3) displayed fairly monotonic velocity changes over most of their monitoring periods. Since double-peaked emitters make up only 3% of all quasars (Strateva et al., 2003), such large monotonic variations are extremely rare––roughly 0.4% of quasars (2/14 of extensively monitored double-peaked emitters 3%). Importantly, none of these objects show a change in sign in their radial velocity curves. Therefore, the velocity changes observed in J0950 cannot be directly attributed to phenomena seen in other quasars.
@@
V.3 A Supermassive Binary Black Hole Scenario for the Radial Velocity Curve
@@
The radial velocity curve of J0950 based on the fourth column of Table 3 is shown in Figure 12. It shows a steady increase of 3500 km s-1 over time; the most striking behavior is that the broad line region of this quasar has undergone radial velocity variations in a manner resembling Keplerian motion in a binary. There certainly are potential caveats, which we discuss in Section V.2. But the radial velocity curve of J0950 so far is consistent with binary motion, which makes it a strong SBHB candidate.
@@
We increased the measurement uncertainties of the radial velocity curve to capture the effects of “jitter” from regular quasar variability (Guo et al. 2019; Doan et al. 2020). Quasars can have perturbations or asymmetries within the broad line region that manifest as small, stochastic radial velocity variations, or jitter, occurring over a timescale of the order of a year or less (e.g., Barth et al. 2015). This subtle effect is directly measurable in two of J0950’s spectra (see Section III.2.1). Doan et al. (2020) empirically determined that jitter can be characterized as fluctuations around a smooth radial velocity curve, distributed approximately according to a Gaussian distribution (see also Guo et al. 2019). The standard deviation of the distribution is on the scale of a few hundred km s-1. Following the methodology from that work, we added a jitter of 200 km s-1 in quadrature to the uncertainties in each radial velocity measurement. The jitter dominates over the measurement errors and is the primary source of uncertainty in the radial velocity at each epoch.
@@
To fit the radial velocity curve shown in Figure 12, we used radvel (Fulton et al. 2018), a package for fitting Keplerian orbits to radial velocity time series. The best-fitting orbital parameters, determined from the radial velocity curve fit in the observed frame, are provided in Table 4. The free parameters in the fitting are the period (), the time of conjunction (), the natural log of the semi-amplitude (ln ), and and , from which the eccentricity () and argument of periapsis () are derived. Given the exploratory nature of fitting the velocity curve of J0950 as a binary and the absence of prior knowledge about the system, we adopted broad, physically motivated priors based on the radial velocity curve. Specifically, we assumed a near-circular orbit as a starting point and adopted Gaussian priors on the period and time of conjunction: a mean period of yr (ensuring that the 1- range does not fall below the 22 yr span of radial velocity measurements available), and a time of conjunction yr. For a circular orbit, the conjunction––the moment when the lower-mass, or secondary BH, passes along the line of sight between the observer and the primary BH––would be expected near that epoch. A hard bound of was applied to the semi-amplitude to restrict unphysical amplitudes, and an eccentricity prior of was imposed. The Markov chain Monte Carlo (MCMC) was run with 20 walkers for 2500 steps each, and the chains converged reliably after steps (Gelman-Rubin statistic 1.01) following an initial burn-in. The resulting posteriors were well-constrained and consistent with the observed signal (see Appendix B for a corner plot), indicating that the priors were non-restrictive.
@@
Since the radial velocity baseline of 22 yr covers only a fraction of the 33-yr orbit, the period, time of conjunction and the semi-amplitude, become partially degenerate. Therefore, the posterior distribution at longer periods is strongly influenced by the assumed prior. For much longer periods, the radial velocity curve would look almost the same over the observed 22 yr window. Likewise, the semi-amplitude can shift slightly depending on the adopted jitter value. These effects are expected for long period systems. In order to improve the sensitivity of the data to longer orbital periods, further monitoring of the radial velocity curve of J0950 is needed for a substantial amount of time.
@@
V.4 Exploration of the Supermassive Binary Black Hole Scenario and Discussion
@@
A lower limit on the total mass of the SBHB in J0950 was initially estimated at M⊙ by Runnoe et al. (2017). We obtained limits on the mass assuming that the bolometric luminosity does not exceed the Eddington luminosity. The bolometric luminosity was estimated by taking the luminosity of J0950’s 2002 spectrum––the brightest state of this object––at 5100 Å, and applying a bolometric correction factor of 10.3 (Richards et al. 2006). This yields a mass of . We estimated another lower mass limit, taking maximum inclination () and a mass ratio of , in the following Keplerian velocity equation,
@@
This equation was derived under the assumption of an unequal mass binary system, in which the less massive, or secondary, BH is active and its BLR emits the observed H line (see arguments in Eracleous et al. 2012; Runnoe et al. 2015). Using all combinations explored by the MCMC algorithm for the period , the velocity semi-amplitude , and eccentricity , the inferred 68%-confidence lower limit on total mass is and the 99% lower limit is . These inferred masses do not necessarily correspond to the lower limits on , , and . All the lower mass limits above are reasonable in the sense that they are less than , a threshold regarded as unphysical. This threshold comes about because the most massive BHs measured are of the order of and because BHs are not expected to accrete material rapidly enough to surpass this mass (e.g., Thomas et al. 2016; Inayoshi and Haiman 2016; McConnell et al. 2012). In line with this argument, the minimum possible inclination for J0950 is . Using the distributions of period and mass, applying Kepler’s law gives 68% and 99% lower limits on the orbital semi-major axis of 0.008 pc and 0.006 pc respectively.
@@
If J0950 is a binary, we attribute the offset broad H emission lines to a BLR associated with the lower mass, or secondary, BH. In this scenario, the tidal interaction with the companion BH may limit the size of the BLR. To explore that possibility, we estimate the “characteristic radius” of the BLR888This is the length that corresponds to the observed time delay, between fluctuations of continuum and a broad line, . using the empirical relation from Bentz et al. (2013) between the characteristic radius and AGN luminosity at 5100 Å,
@@
with and . We applied this relation even though it is appropriate only for a single BH and does not account for the dynamical effects in a binary system to get a characteristic radius of the BLR of pc. This value is somewhat higher than our inferred lower limit on the binary semi-major axis of pc. The extent of the line-emitting region in the BLR is possibly 45 times greater than (see discussion in Runnoe et al. 2015). Since we only have a lower limit on the orbital semi-major axis, there appears to be no conflict between this separation and the characteristic radius of the BLR.
@@
If J0950 does indeed harbor a binary, extrapolating from the fit to its radial velocity curve, it should start turning over soon. It is therefore critical that we observe this object in the near future, at a rate of at least one spectrum per year to see if this prediction is borne out. It is also important to note that our implementation of radial velocity jitter discussed in Section III.5 is simple, and more sophisticated prescriptions of jitter should be developed and employed. The description of jitter would impact the uncertainties on the orbital parameters that are inferred from fitting a radial velocity curve. Regardless of whether or not J0950 is a binary, it still displays an exceptional behavior that warrants detailed study.
V.5 On the Detectability of Doppler Boosting in the Light Curves